# Muffin: Mitigating Unhelpfulness in Emotional Support Conversations with Multifaceted AI Feedback

### v2

@@ Abstract
Emotional support conversation systems are designed to alleviate users’ emotional distress and assist them in overcoming their challenges. While previous studies have made progress, their models occasionally generate unhelpful responses, which are intended to be supportive but instead have counterproductive effects. Since unhelpful responses can hinder the effectiveness of emotional support, it is crucial to mitigate them within conversations. Our solution is motivated by two principal considerations: (1) multiple facets of emotional support are expected to be considered when developing emotional support conversation models, and (2) directly reducing the probability of generating unhelpful responses can effectively mitigate their occurrence. Accordingly, we introduce a novel model-agnostic framework named Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). It first employs a multifaceted AI feedback module designed to assess the helpfulness model responses across various facets of emotional support. Leveraging contrastive learning, Muffin then reduces the unhelpful responses’ likelihoods. To validate the effectiveness of our proposed framework, we apply Muffin to various previous emotional support generation models, including the state-of-the-art. Experimental results demonstrate that Muffin can significantly mitigate unhelpful response generation while enhancing response fluency and relevance.
@@
Jiashuo WANG, Chunpu XU, Chak Tou LEONG, Wenjie LI, Jing LI Hong Kong Polytechnic University {csjwang,cswjli,jing-amelia.li}@comp.polyu.edu.hk {chun-pu.xu,chak-tou.leong}@connect.polyu.hk
@@ 1 Introduction
Emotional support conversation systems (supporters) are designed to generate responses that can buffer the emotional distress experienced by users (help-seekers) and help users to work through the challenges they are confronting Liu et al. (2021). Recently, many studies have contributed to this field Deng et al. (2023); Zhao et al. (2023); Cheng et al. (2022); Tu et al. (2022). Despite some success, their models occasionally generate well-intended responses that produce a counterproductive support effect, i.e., exacerbating the negative emotional states of users or inhibiting effective problem-solving, as shown in Figure 1. In the psychology and communication theories, these failed support attempts are termed “unhelpful messages” Greene and Burleson (2003); Burleson (1985).
@@
Accordingly, this paper introduces a novel, model-agnostic framework called Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). For D1, we design a multifaceted AI feedback module within the Muffin framework. This module assesses whether a specific response is unhelpful from multiple facets of emotional support. Leveraging the advanced capabilities of recent large language models (LLMs), we implement this module by instruction-tuning LLaMA, avoiding inefficient and expensive human feedback collection. Then, we continue optimizing an emotional support conversation model with its previous training objective and a new one to implement D2. The additional objective is to minimize the likelihood of unhelpful responses, which is implemented by contrasting unhelpful responses, identified by the feedback module, and the other (non-unhelpful) ones. Through these two steps, we aim to mitigate the unhelpful responses generated by a given emotional support conversation model. Experimental results highlight the effectiveness of our framework, demonstrating that Muffin can enhance the helpfulness of previous emotional support conversation models, including those recognized as state-of-the-art. The main contributions of this work are:
@@ 2.1 Emotional Support Conversation
In the domain of emotional support conversation generation, prior studies have achieved some success. Specifically, they have each emphasized and incorporated different facets of emotional support, such as empathetic expression, communication skills, and response coherence, into their respective models. Some of them consider a single facet. For example, MultiESC Cheng et al. (2022) only considers the communication skill efficacy of responses by planning response strategies. KEMI Deng et al. (2023) incorporates related external knowledge for response generation. Although the response coherence is enhanced, the efficacy of communication skills is ignored. Li et al. (2024) enhance the empathetic expression of LLMs through chain-of-though. Beyond these, MISC Tu et al. (2022) generates supportive responses considering both commonsense and communication skills, thereby enhancing two facets of emotional support. TranESC Zhao et al. (2023) incorporate commonsense, communication skills, and emotional elements into response generation. Most all these models are optimized by minimizing the negative log-likelihood of golden responses, instead of the unhelpful response likelihood.
@@ 2.2 Contrastive Learning for Text Generation
Contrastive learning was initially employed to learn meaningful representations by contrasting positive and negative samples in the field of natural language processing Logeswaran and Lee (2018); Gutmann and Hyvärinen (2012). Recently, it has been applied to text generation Jiashuo et al. (2023); Zheng et al. (2023); Liu et al. (2022), achieving impressive performance across various settings. During training, the model is exposed to a range of “hard” negative and positive samples through contrastive learning, enabling the model to distinguish preferred outputs from less desirable ones. Consequently, selecting positive and negative samples is crucial in this process. In this paper, responses generated by the model that are deemed unhelpful are negative samples, while all other responses are considered positive samples.
@@ 3.2 Problem Definition
Our primary goal is to mitigate the generation of unhelpful responses. Rather than training a new model from scratch, we aim to refine a pre-trained emotional support conversation model with the dataset it was originally trained on. This process unfolds as follows. Let represent the model trained on the dataset , where denotes the model parameters. Each instance in is denoted as , with as the input and as the expected output. Usually, is the conversation context, but it contains additional related information in some models. Assume that there are samples with labels . These samples are the diverse beam search generation results of . As for the label , it represents feedback to indicate whether the sample is unhelpful () or not unhelpful (). Our objective is to refine the model’s parameters such that the likelihood of generating unhelpful samples is reduced relative to helpful ones. In this process, we only modify the training process, ensuring that the model’s architecture and the inference mechanism remain untouched. Moreover, our approach is model-agnostic. This implies that can be any deep learning model designed and trained for emotional support conversations.
@@ 4 Method
The overall framework of Muffin is outlined in Figure 2. It is composed of two principal components, each specifically designed for deliberations in Section 1: P1: Consideration of Multiple Facets and P2: Direct Mimimization of Unhelpful Response Probality, respectively. The multifaceted AI feedback module aims to identify whether a response from multiple facets of emotional support is unhelpful. The unhelpfulness mitigation module mitigates the likelihood of unhelpful responses by contrasting helpful and unhelpful responses.
@@ 5.1 Experimental Setups
The ESConv dataset Liu et al. (2021) is used to train the aforementioned base models. ESConv is a benchmark for emotional support conversations, comprising approximately 1K conversations with 13K utterances. All base models, with the exception of TransESC, follow the original division of ESConv for training, validation, and testing, using an 8:1:1 ratio. TransESC employs a random split while maintaining the same ratio. Notably, each model adopts different data preprocessing methods. We adhere to each base model’s specific data division and pre-processing.
@@ 5.2 Facets of Emotional Support
Empathetic expressions signify the supporter’s interest in and comprehension of the help-seeker’s perspective. Conversely, their absence can impede conversation engagement and obstruct establishing a trust-based relationship between the supporter and the help-seeker Morse et al. (1992). While empathy encompasses various aspects Paiva et al. (2017); Batson (2009); Bohart and Greenberg (1997), we adopt the comprehensive framework proposed by Sharma et al. (2020). It identifies three empathy communication mechanisms, i.e., emotional reaction, interpretations, and explorations, each assessed across three levels: no communication, weak communication, and strong communication. Our work considers responses exhibiting no empathy across all mechanisms unhelpful. For example, “sleeplessness can result in upsetness,” which inappropriately offers mere information, is considered unhelpful in empathetic expression, especially when responding to a statement like “I am upset”.
@@ 5.4 Emotional Support Response Generation
Following previous work Deng et al. (2023); Liu et al. (2021), we compare the base model and its corresponding Muffin model on five aspects, which are (1) Fluency: which model’s response is more fluent? (2) Identification: which model’s response is more skillful in identifying the user’s problem? (3) Comforting: which model’s response is better at comforting the user? (4) Suggestion: which model can give more helpful and informative suggestions? (5) Overall: which model’s response is generally more helpful from the aspect of the help-seeker? Specifically, for each -Muffin pair, we randomly select 100 instances for comparison. Then, we ask four unique human evaluators to vote which response is better. They can select “tie” if both responses are considered equal. We average their results as the final result.
@@
Figure 5 summarizes the A/B test results on BlenderBot-Joint Liu et al. (2021), KEMI Deng et al. (2023), and MultiESC Cheng et al. (2022), along with their corresponding Muffin models. These three settings are selected for their significant performance in automatic evaluation. The inter-rater agreement, i.e., Fleiss’ Kappa Fleiss (1971), is , implying fair agreement. Our Muffin models are regarded as more helpful in general. Responses generated by Muffin models are slightly more fluent than those generated by base models. Compared with the corresponding base model , the Muffin model shows more powerful capability in identifying the help-seeker’s problem. Moreover, Muffin models can generate responses that have better effects to comfort the users than . Annotators also prefer responses generated by Muffin because of their more helpful and informative suggestions. These results prove that Muffin indeed mitigates the unhelpfulness of emotional support conversation models.
Model | B-1 | B-2 | B-3 | B-4 | R-L | METEOR | CIDEr | Extreme |
Joint | 18.77 | 7.54 | 3.79 | 2.15 | 17.72 | 7.59 | 17.38 | 50.96 |
Muffin | 20.59 | 8.38 | 4.26 | 2.54 | 18.35 | 8.18 | 19.12 | 51.46 |
Muffin (emp.) | 19.58 | 8.06 | 4.04 | 2.33 | 18.46 | 7.87 | 19.61 | 51.43 |
Muffin (skill) | 18.68 | 7.51 | 3.81 | 2.19 | 17.98 | 7.67 | 17.84 | 50.98 |
Muffin (cohr.) | 20.04 | 8.10 | 4.04 | 2.26 | 18.24 | 7.95 | 18.07 | 51.31 |
Ablation study
The findings, as presented in Table 2, reveal that relying solely on AI feedback from a single facet for unhelpfulness mitigation results in diminished performance in automatic evaluations when compared to the comprehensive Muffin model. This underscores that considering multiple facets of emotional support when building emotional support conversation models (P1) is necessary. Another insight from Table 2 is that all three ablated models outperform the base model. It suggests that the proposed solution, directly mitigating unhelpful responses (P2), is reasonable and effective. Moreover, the results also indicate that the quality of helpful and unhelpful responses will influence the effects of unhelpfulness mitigation.
Conversation History | |
Help-seeker : I’m an alcoholic. I get that thought in my head and its like I become obsessed. |
Supporter : Why not try to put attention on some other things instead of a beer? |
Help-seeker : Hmmm like what? |
Model Generated Responses | |
Joint |
You can try to be normal. That is a great idea. (Lacking empathy) |
Muffin |
Yes, there are some things that you can do, but you have to be aware of your own mental health, that’s why it’s important to know you are doing the right thing. |
MultiESC |
It sounds like you have a lot on your plate and you want to be normal. (Inconsistent) I can understand how that would make you feel. |
Muffin |
I think you should focus on yourself instead of a beer. Do you have any hobbies that you enjoy? |
KEMI |
I think you should try to find some new things to do to relax. |
Muffin |
Have you tried meditation? It can help you relax and relax. |
Case study
6 Conclusion
Limitations
Ethical Considerations
Appendix A Multifacetd AI Feedback Module
A.1 Multiple Facets of Emotional Support
Facet |
Context |
Unhelpful Response |
Empathetic Expression |
Help-seeker: Just wanna sleep. Can’t motivate at all. I don’t know what to do. |
Supporter: As Nas says, sleep is the cousin of death. It’s the kindest form of oblivion. |
Help-seeker: What do you want to be when you grow up? I don’t know what to do with my life or what I truly love when all I really want is not to exist. |
Supporter: But what does fully grown up mean? I have my own house and a job etc. Am I fully grown up now? |
Communication Skill Efficiency |
Help-Seeker: I do nothing but think about how I’m worthless and how I shouldn’t be here. I’ve never tried or contemplated suicide. I’ve always wanted to fix my issues, but I never get around to it. How can I change my feeling of being worthless to everyone? |
Supporter: Still, you should watch out for this because it can be a sign of worsening depression. |
Help-Seeker: I try to have a “heart-to-heart” with my mom, but it usually winds up being messy and frustrating afterward. Thank you for your advice and words of encouragement. Sorry if that sounds silly. |
Supporter: Stop being so afraid…you really have nothing to lose :) Well, you are right. |
Response Coherence |
Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And I am just always tired of everything these days. |
Supporter: Could you please explain why you are enjoying yourself? |
Supporter: Yes, what happened? Why are you tired? Please can you explain? Everything going ok? |
Help-seeker: I think it’s my depression. I have a very hard time getting motivated to do anything. Nothing seems worth it. |
Help-seeker: I am upset. |
Supporter: Everyone can be happy. Can you tell me more? |
Empathetic Expression
Communication Skill Efficiency
Response Coherence
Context | |
Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And, I am just always tired, tired of everything these days. |
Original Response |
Yes, What happened? Why are you tired? Please can you explain? |
Incoherent Responses |
I am glad to hear you are feeling a little better! Yes, bad management is so toxic. Even with great coworkers in a job you love, horrible management can ruin it quickly. (From another dialogue) |
A.2 Model Tuning and Module Performance
We instruction-tuned LLaMA111https://huggingface.co/decapoda-research/llama-7b-hf to equip the model with the capability required for unhelpful response classification tasks.222The implementation referred to is available at https://huggingface.co/decapoda-research/llama-7b-hf Furthermore, we initialized the LoRA weights using a low-rank adapter that was fine-tuned on the Stanford Alpaca dataset.333https://huggingface.co/tloen/alpaca-lora-7b We evaluated the capabilities of GPT-3.5 and GPT-4 by invoking the gpt-3.5-turbo and gpt-4 models, respectively, through their API444https://platform.openai.com/docs/api-reference/chat/create during the period from December 2023 to January 2024. We set the temperature parameter to to ensure deterministic output generation. The detailed results of this evaluation are detailed in Table 6.
Model |
GPT-3.5 |
GPT-4 |
LLaMA (Vanilla) |
LLaMA (Tuned) |
Empathetic Expression | accuracy |
81.88 |
69.58 |
19.09 |
83.50 |
F1 Score |
54.89 |
61.58 |
16.03 |
74.12 |
Skill Efficiency | accuracy |
69.56 |
79.01 |
84.96 |
90.43 |
F1 Score |
53.22 |
66.14 |
45.93 |
81.83 |
Response Coherence | accuracy |
33.47 |
78.76 |
33.31 |
92.98 |
F1 Score |
25.26 |
72.62 |
24.98 |
92.06 |
Aggregated Feedback | accuracy |
57.11 |
78.01 |
59.08 |
90.72 |
F1 Score |
41.99 |
77.19 |
40.96 |
89.86 |
We also conduct human evaluation to examine the module. In particular, we randomly sample responses from the ESConv dataset and model-generated responses. For each response, an annotator was asked to assess whether it is unhelpful. We compared the annotations with the multifaceted AI feedback, and the consistency rate was . We found that the multifaceted AI module is stricter than human annotators. It is evident by the fact that for instances which the multifaceted AI feedback and human annotations are different, the multifaceted AI feedback tends to consider the response unhelpful.
Appendix B Experimental Setup
B.1 Devices and Environment
We employed Pytorch555https://pytorch.org/ for the implementation of all models, and conducted our experiments using Nvidia GeForce RTX 3090 GPUs.
B.2 ESConv Dataset
B.3 Base Models
Vanilla
is a vanilla BlenderBot Roller et al. (2021) trained on the dataset ESConv. We used the small version666https://huggingface.co/facebook/blenderbot_small-90M of BlenderBot in experiments following previous studies.
Joint
is built upon the backbone of BlenderBot and is specially trained to generate responses along with an expected communication skill at the beginning of each response Liu et al. (2021)777https://github.com/thu-coai/Emotional-Support-Conversation. The focused facet of this model is skill efficiency.
MultiESC
Cheng et al. (2022)888https://github.com/lwgkzl/multiesc is an emotional support conversation model that mainly focuses on the communication skill efficiency. It predicts the strategies (communication skills) in the next several turns considering the the user’s state.
KEMI
incorporates various knowledge for a mixed-initiative conversation model, which can provide emotional support Deng et al. (2023)999https://github.com/dengyang17/KEMI. In this process, the response coherence is also enhanced.
TransESC
Zhao et al. (2023)101010https://github.com/circle-hit/TransESC predicts the transitions of the user’s emotion, the communication skills, and the conversation keywords. Then, such information is used for response generation. This model takes into account more than one facet of emotional support; however, it is not optimized to reduce the likelihood of unhelpful responses like other models.
B.4 Human Evaluation
Initially, the annotators were briefed about the nature of emotional support conversations to ensure a comprehensive understanding of the task’s objectives. During the rating process, they were suggested to imagine themselves as the help-seekers within the conversations. After being provided with the conversation context, the annotators then proceeded to compare two generated responses, shown as LABEL:fig:huam_rating. Moreover, we prioritized the comfort and well-being of our annotators, advising them to pause or cease the annotation process if they encountered any content that made them feel uncomfortable. Annotators were paid at a rate of 1.52 times their local hourly minimum wage. Based on annotators’ feedback, it was estimated that approximately 40 seconds were spent on evaluating each response pair.
Appendix C Response Generation Model Performance
C.1 Multifaceted AI Feedback
In addition to presenting the primary results and conducting an ablation study, we provide the multifaceted AI feedback results as a reference, as outlined in Table 7. For each model, we calculate the frequency of generating unhelpful responses on the ESConv testing set. Specifically, we utilize the multifaceted AI feedback module to identify the helpful (non-unhelpful) responses and compute their percentage, displayed in the left subtable. Furthermore, we analyze the helpful response percentage when each model generates ten responses using diverse beam search, reported in the right subtable. It’s worth noting that the assertion “approximately 30% of responses generated by MultiESC on the ESConv benchmark are identified as unhelpful” is derived from the findings presented in this table.

### v3

@@ Abstract
Emotional support conversation systems are designed to alleviate users’ emotional distress and assist them in overcoming their challenges. While previous studies have made progress, their models occasionally generate unhelpful responses, which are intended to be supportive but instead have counterproductive effects. Since unhelpful responses can hinder the effectiveness of emotional support, it is crucial to mitigate them within conversations. Our solution is motivated by two principal considerations: (1) multiple facets of emotional support are expected to be considered when developing emotional support conversation models, and (2) directly reducing the probability of generating unhelpful responses can effectively mitigate their occurrence. Accordingly, we introduce a novel model-agnostic framework named Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). It first employs a multifaceted AI feedback module designed to assess the helpfulness model responses across various facets of emotional support. Leveraging contrastive learning, Muffin then reduces the unhelpful responses’ likelihoods. To validate the effectiveness of our proposed framework, we apply Muffin to various previous emotional support generation models, including the state-of-the-art. Experimental results demonstrate that Muffin can significantly mitigate unhelpful response generation while enhancing response fluency and relevance. We release our codes at https://github.com/wangjs9/Muffin.
@@
Jiashuo Wang, Chunpu Xu, Chak Tou Leong, Wenjie Li††thanks: Corresponding author., Jing Li Hong Kong Polytechnic University {csjwang,cswjli,jing-amelia.li}@comp.polyu.edu.hk {chun-pu.xu,chak-tou.leong}@connect.polyu.hk
@@ 1 Introduction
Emotional support conversation systems (supporters) are designed to generate responses that can buffer the emotional distress experienced by users (help-seekers) and help users to work through the challenges they are confronting Liu et al. (2021). Recently, many studies have contributed to this field Deng et al. (2023); Zhao et al. (2023); Cheng et al. (2022); Tu et al. (2022). Despite great success, their models occasionally generate well-intended responses that produce a counterproductive support effect, i.e., exacerbating the negative emotional states of users or inhibiting effective problem-solving, as shown in Figure 1. In the psychology and communication theories, these failed support attempts are termed “unhelpful messages” Greene and Burleson (2003); Burleson (1985).
@@
Accordingly, this paper introduces a novel model-agnostic framework called Mitigating unhelpfulness with multifaceted AI feedback for emotional support (Muffin). For D1, we design a multifaceted AI feedback module within the Muffin framework. This module assesses whether a specific response is unhelpful from multiple facets of emotional support. Leveraging the advanced capabilities of recent large language models (LLMs), we implement this module by instruction-tuning LLaMA, avoiding inefficient and expensive human feedback collection. Then, we continue optimizing an emotional support conversation model with its previous training objective and a new one to implement D2. The additional objective is to minimize the likelihood of unhelpful responses, which is implemented by contrasting unhelpful responses, identified by the feedback module, and the other (non-unhelpful) ones. Through these two steps, we aim to mitigate the unhelpful responses generated by a given emotional support conversation model. Experimental results highlight the effectiveness of our framework, demonstrating that Muffin can enhance the helpfulness of previous emotional support conversation models, including those recognized as state-of-the-art. The main contributions of this work are:
@@ 2.1 Emotional Support Conversation
In the domain of emotional support conversation generation, prior studies have achieved some success. Specifically, they have each emphasized and incorporated different facets of emotional support, such as empathetic expression, communication skills, and response coherence, into their respective models. Some of them consider a single facet. For example, MultiESC Cheng et al. (2022) only considers the communication skill efficacy of responses by planning response strategies. KEMI Deng et al. (2023) incorporates related external knowledge for response generation. Although the response coherence is enhanced, the efficacy of communication skills is ignored. Li et al. (2024) enhance the empathetic expression of LLMs through chain-of-though. Beyond these, MISC Tu et al. (2022) generates supportive responses considering both commonsense and communication skills, thereby enhancing two facets of emotional support. TranESC Zhao et al. (2023) incorporates commonsense, communication skills, and emotional elements into response generation. Most all these models are optimized by minimizing the negative log-likelihood of golden responses, instead of the unhelpful response likelihood.
@@ 2.2 Contrastive Learning for Text Generation
Contrastive learning was initially employed to learn meaningful representations by contrasting positive and negative samples in the field of natural language processing Logeswaran and Lee (2018); Gutmann and Hyvärinen (2012). Recently, it has been applied to text generation Wang et al. (2024); Zheng et al. (2023); Liu et al. (2022), achieving impressive performance across various settings. During training, the model is exposed to a range of “hard” negative and positive samples through contrastive learning, enabling the model to distinguish preferred outputs from less desirable ones. Consequently, selecting positive and negative samples is crucial in this process. In this paper, responses generated by the model that are deemed unhelpful are negative samples, while all other responses are considered positive samples.
@@ 3.2 Problem Definition
Our primary goal is to mitigate the generation of unhelpful responses. Rather than training a new model from scratch, we aim to refine a pre-trained emotional support conversation model with the dataset it was originally trained on. This process unfolds as follows. Let represent the model trained on a dataset , where denotes the model parameters. Each instance in is denoted as , with as the input and as the expected output. Usually, is the conversation context, but it contains additional related information in some models. Assume that there are samples with labels . These samples are the diverse beam search generation results of . As for the label , it represents feedback to indicate whether the sample is unhelpful () or not unhelpful (). Our objective is to refine the model’s parameters such that the likelihood of generating unhelpful samples is reduced relative to helpful ones. In this process, we only modify the training process, ensuring that the model’s architecture and the inference mechanism remain untouched. Moreover, our approach is model-agnostic. This implies that can be any deep learning model designed and trained for emotional support conversations.
@@ 4 Method
The overall framework of Muffin is outlined in Figure 2. It is composed of two principal components, each specifically designed for deliberations in Section 1: D1: Consideration of Multiple Facets and D2: Direct Mimimization of Unhelpful Response Probality, respectively. The multifaceted AI feedback module aims to identify whether a response from multiple facets of emotional support is unhelpful. The unhelpfulness mitigation module mitigates the likelihood of unhelpful responses by contrasting helpful and unhelpful responses.
@@ 5.1 Experimental Setups
The ESConv dataset Liu et al. (2021) is used to train the aforementioned base models. ESConv is a benchmark for emotional support conversations, comprising approximately 1K conversations with 31K utterances. All base models, with the exception of TransESC, follow the original division of ESConv for training, validation, and testing, using an 8:1:1 ratio. TransESC employs a random split while maintaining the same ratio. Notably, each model adopts different data preprocessing methods. We adhere to each base model’s specific data division and pre-processing.
@@ 5.2 Facets of Emotional Support
Empathetic expressions signify the supporter’s interest in and comprehension of the help-seeker’s perspective. Conversely, their absence can impede conversation engagement and obstruct establishing a trust-based relationship between the supporter and the help-seeker Morse et al. (1992). While empathy encompasses various aspects Wang et al. (2021); Paiva et al. (2017); Bohart and Greenberg (1997), we adopt the comprehensive framework proposed by Sharma et al. (2020). It identifies three empathy communication mechanisms, i.e., emotional reaction, interpretations, and explorations, each assessed across three levels: no communication, weak communication, and strong communication. Our work considers responses exhibiting no empathy across all mechanisms unhelpful. For example, “sleeplessness can result in upsetness,” which inappropriately offers mere information, is considered unhelpful in empathetic expression, especially when responding to a statement like “I am upset”.
@@ 5.4 Emotional Support Response Generation
Following previous work Deng et al. (2023); Liu et al. (2021), we compare the base model and its corresponding Muffin model on five aspects, which are (1) Fluency: which model’s response is more fluent? (2) Identification: which model’s response is more skillful in identifying the user’s problem? (3) Comforting: which model’s response is better at comforting the user? (4) Suggestion: which model can give more helpful and informative suggestions? (5) Helpfulness: which model’s response is generally more helpful from the aspect of the help-seeker? Specifically, for each -Muffin pair, we randomly select 100 instances for comparison. Then, we ask four unique human evaluators to vote which response is better. They can select “tie” if responses are considered equal. We average their results as the final result.
@@
Figure 5 summarizes the A/B test results on BlenderBot-Joint Liu et al. (2021), KEMI Deng et al. (2023), and MultiESC Cheng et al. (2022), along with their corresponding Muffin models. These three settings are selected for their significant performance in automatic evaluation. The inter-rater agreement, i.e., Fleiss’ Kappa Fleiss (1971), is , implying fair agreement. Our Muffin models are regarded as more helpful in general, as evidenced by their higher Helpfulness. Responses generated by Muffin models are slightly more fluent than those generated by base models. We also observe that ’ties’ are common in evaluations of response fluency, mainly because the responses generated are typically fluent. Compared with the corresponding base model , the Muffin model shows some more powerful capability in identifying the help-seeker’s problem. Moreover, Muffin models can generate responses that have better effects to comfort the users than . Annotators also prefer responses generated by Muffin because of their more helpful and informative suggestions. These results prove that Muffin indeed mitigates the unhelpfulness of emotional support conversation models.
@@
Model | B-1 | B-2 | B-3 | B-4 | R-L | METEOR | CIDEr | Extreme |
Joint | 18.77 | 7.54 | 3.79 | 2.15 | 17.72 | 7.59 | 17.38 | 50.96 |
Muffin | 20.59 | 8.38 | 4.26 | 2.54 | 18.35 | 8.18 | 19.12 | 51.46 |
Muffin (emp.) | 19.58 | 8.06 | 4.04 | 2.33 | 18.46 | 7.87 | 19.61 | 51.43 |
Muffin (skill) | 18.68 | 7.51 | 3.81 | 2.19 | 17.98 | 7.67 | 17.84 | 50.98 |
Muffin (cohr.) | 20.04 | 8.10 | 4.04 | 2.26 | 18.24 | 7.95 | 18.07 | 51.31 |
Helpfulness Evaluation via AI Feedback
We provide the multifaceted AI feedback results as a reference, as outlined in Table 2. We utilize the multifaceted AI feedback module to identify various models’ helpful (non-unhelpful) responses and compute their percentage, displayed in the left subtable. Furthermore, we analyze the helpful response percentage when each model generates ten responses using diverse beam search, reported in the right subtable.111The assertion “approximately 30% of responses generated by MultiESC on the ESConv benchmark are identified as unhelpful” is derived from the findings presented in this table.